OpenStudy (anonymous):

Consider two points A (a, o) and B (-a, 0) belongs to R*. Determining (Ln) the set of points M (x, y) such that: cos (vecteurMA, vecteurMA) = (root 3) / 2

OpenStudy (anonymous):

cos (vecteurMA,vecteurMB) sorry not cos ( vecteur MA,vecteur MA) So cos (vecteurMA,vecteurMB)=cos(pi/6)=MA*MB/vecteurMA.vecteurMB What else ?

OpenStudy (loser66):

We have M (x,y) , A (a,0) and B (-a,0) \[cos (\vec{MA}, \vec {MB})= \frac{<MA,MB>}{|MA||MB|}=\frac{\sqrt3}{2}\] Calculate the middle term: MA = (x - a, y) MB = (x + a, y) |MA| = \(\sqrt{(x-a)^2 +y^2}\) |MB|= \(\sqrt{(x+a)^2+y^2}\) replace all \[\frac{<((x-a),y),((x+a),y)>}{\sqrt{(x-a)^2 +y^2}\sqrt{(x+a)^2+y^2}}=\frac{\sqrt3}{2}\] \[\rightarrow 2(x^2-a^2+y^2)=\sqrt{3}\sqrt{(x-a)^2 +y^2}\sqrt{(x+a)^2+y^2}\] now, square both sides and simplify to consider what the form it is. hehehe... We have an outlet there. Now, your turn.